a park is in the shape of a rectangle 4 km long and 3km wide. you are walking from point a to point b how much shorter is your walk diagonaly across the park rather than along of its 2 sides?
Answer by richwmiller(17219)

Don't know where point a is nor point b
So can't tell any of the distances.
now if a and b are on the diagonal corners of the park
we can add 4+3=7
16+9=25
sqrt(25)=5
7-5=2 km difference